# Assessing Large Language Models in Updating Their Forecasts with New Information

### v1

@@ Abstract
Prior work has largely treated future event prediction as a static task, failing to consider how forecasts and the confidence in them should evolve as new evidence emerges. To address this gap, we introduce EvolveCast, a framework for evaluating whether large language models appropriately revise their predictions in response to new information. In particular, EvolveCast assesses whether LLMs adjust their forecasts when presented with information released after their training cutoff. We use human forecasters as a comparative reference to analyze prediction shifts and confidence calibration under updated contexts. While LLMs demonstrate some responsiveness to new information, their updates are often inconsistent or overly conservative. We further find that neither verbalized nor logits-based confidence estimates consistently outperform the other, and both remain far from the human reference standard. Across settings, models tend to express conservative bias, underscoring the need for more robust approaches to belief updating.
@@ 1 Introduction
Large language models (LLMs) have achieved strong performance across a wide range of NLP tasks, particularly those involving factual recall and reasoning over known information (Lin et al., 2022; Guo et al., 2023) including temporal information (Ding et al., 2025). However, most evaluations remain static and retrospective: they assess what models know about the world up to their training cutoff. In contrast, real-world decision-making often requires reasoning about uncertain future events, where outcomes are not yet known and relevant information evolves over time.
Forecasting plays a central role in domains such as policymaking, science, and technology. It requires anticipating future developments based on incomplete or uncertain evidence, and adjusting those beliefs as new signals arrive. Unlike fact retrieval tasks, forecasting is inherently temporal and dynamic. For instance, given the question, “Will GPT-5 be released in 2025?”, an LLM might answer “Yes” based on existing trends in AI development. However, without a way to incorporate emerging information, such as announcements or delays, such forecasts remain static, and may quickly become outdated or misleading. To illustrate the dynamic nature of forecasting, see the example in Figure 1.
Prior work has begun to explore the forecasting capabilities of LLMs (Jin et al., 2020; Zou et al., 2022; Yuan et al., 2024a). These efforts typically frame forecasting as a static problem, where the model produces a categorical answer based on its existing knowledge. Some studies incorporate external information, including textual evidence such as news (Ye et al., 2024; Halawi et al., 2024), but use this information to improve the model’s final prediction, not to evaluate how forecasts change over time. More recently, forecasting has been reframed as a probabilistic task, where models are expected to produce confidence scores in addition to predictions (Karger et al., 2024; Yuan et al., 2025). In this setting, the goal is not just to answer the mentioned question with “Yes”, but to provide a confidence level, e.g., 70%, that reflects the model’s uncertainty. These works though showed that even when models make correct predictions, their confidence is often miscalibrated.
Yet even confidence-aware predictions remain static if models cannot revise their beliefs. In practice, forecasters routinely update their views as new information becomes available. For example, if a forecaster initially assigns a 55% chance to the release of GPT-5 in 2025, but later encounters a credible report stating “Sam Altman has hinted at a major announcement next week,” they might revise their estimate to 70%. A capable forecasting model should exhibit similar adaptive behavior, not only producing reasonable confidence scores, but updating them in response to emerging signals. Evaluating belief dynamics provides insight into how well models reason under uncertainty and adapt over time, which remains a blind spot in most LLM assessments.
This paper introduces EvolveCast, a framework for evaluating whether language models revise their forecasts when presented with new information. Rather than focusing on static predictions, we assess whether models adjust both their predicted outcome and associated confidence in response to emerging evidence. To this end, we construct a benchmark of forecasting scenarios where information relevant to the forecast becomes available after a model’s training cutoff. We evaluate LLM predictions with and without access to this new information, and evaluate the resulting belief changes against those made by human forecasters under similar conditions. In an additional setting, we also provide the model with a time series of historical human forecasts up to the moment of the news, allowing it to optionally leverage trends in human belief as a contextual reference without treating it as ground truth. This enables a softer form of conditioning that mimics the information a real-world forecaster might observe, without prescribing correct answers.
Model updating its forecast in the correct direction but wrong magnitude.QuestionModelReference Q: Will GPT-5 be released in 2025? Initial forecast (t=0t{=}0) Prediction: Yes
Confidence: $p_{0}=55\%$
News update: “Sam Altman hints at major 2025 release.” Updated forecast (t=1t{=}1) Prediction: Yes
Confidence: $p_{t}=60\%$
Direction: Up Magnitude: $5\%$
@@
Model under-reacts vs. reference: 155=1015{-}5=10 $p$ shortfall
Figure 1: Illustrative example of belief updating in EvolveCast. A forecasting question asks: “Will GPT-5 be released in 2025?”. Both the model and the human reference start at 55% confidence. After a news update (“Sam Altman hints at major 2025 release”), the human reference revises upward to 70% (an increase of 15%), while the model increases only to 60% (an increase of 5%). The model captures the correct direction of change but underestimates its magnitude, resulting in a 10% shortfall.
@@
Our evaluation focuses on the direction and magnitude of prediction shifts and the calibration of updated confidence scores as exampled in Figure 1. Our findings show that while language models can adjust their forecasts in response to new information, their updates remain limited: belief revisions are often conservative, confidence estimates vary by method without a clear advantage, and both fall short of human reference forecasts. These challenges highlight fundamental difficulties in modeling belief dynamics and underscore the need for evaluation frameworks that extend beyond static forecasting accuracy. EvolveCast provides such a framework, offering a principled setting for analyzing how models adapt their predictions as the world evolves.
@@ 2 Related Work
A growing body of research has investigated the forecasting capabilities of language models, though most efforts emphasize event prediction rather than belief dynamics or confidence calibration. OpenForecast (Wang et al., 2025) introduces a large-scale benchmark for open-ended multi-step forecasting, but focuses primarily on model accuracy rather than the calibration of confidence estimates. ForecastBench (Karger et al., 2024) frames forecasting as a probabilistic task with evolving predictions over time, but does not explicitly examine how models adjust their beliefs in light of new information. Time-R1 (Liu et al., 2025) presents a reinforcement learning framework to endow smaller LLMs with temporal understanding and future event generation capabilities, showing strong performance on forecasting tasks beyond training cutoff. While our focus also involves forecasting under temporal uncertainty, unlike previous we evaluate belief revisions as new signals emerge. Beyond forecasting, several benchmarks target reasoning and plausibility assessment. COPA (Roemmele et al., 2011) and HellaSwag (Zellers et al., 2019) evaluate causal and commonsense reasoning via multiple-choice inference tasks. PRobELM (Yuan et al., 2024b) examines models’ ability to rank outcomes by plausibility, drawing on general world knowledge. While these evaluations probe important reasoning capabilities, they do not test how models reason under uncertainty or update beliefs over time.
@@ 3.1 Task Definition
We formalize EvolveCast as a dynamic forecasting task, where the goal is to evaluate whether a language model revises its predictions and confidence in response to new information in a manner comparable to a reference standard. Let $q$ denote a forecasting question about a future event (e.g., “Will GPT-5 be released in 2025?”). At time $t$, new information $x_{t}$ (e.g., a news headline or update) becomes available, and we measure how the model’s belief changes when conditioned on it. For binary questions ($\mathcal{Y}=\{\text{Yes},\text{No}\}$), the model’s belief at time $t$ reduces to a scalar confidence $p_{t}=P_{t}(\text{Yes}\mid x_{t})\in[0,1]$, and the belief update is defined as $\Delta p=p_{t}-p_{0}$.
@@ 3.2 Evaluation Criteria
We evaluate the quality of model belief updates by comparing the change in model confidence, $\Delta p=p_{t}-p_{0}$, to the corresponding reference shift $\Delta h=h_{t}-h_{0}$, on three complementary aspects:
@@
We use Mean Directional Accuracy (MDA) to assess whether the model and reference forecasts tend to update their beliefs in the same direction across examples. For a set of $N$ instances, MDA is defined as:
@@
where $\Delta p_{i}$ and $\Delta h_{i}$ are the model and reference forecast deltas for the $i$-th instance. We apply a minimal change threshold $\epsilon$ to avoid counting negligible shifts as meaningful updates. Specifically, we treat both $\Delta p_{i}$ and $\Delta h_{i}$ as zero when $|\Delta|<\epsilon$, and consider the signs to match in such cases. Let $\tilde{\Delta}_{p}$ denote the thresholded version of $\Delta p$, where $\tilde{\Delta}_{p}=0$ if $|\Delta p|<\epsilon$, and similarly for $\Delta h$. Precision, Recall, and F1 (micro-averaged) are also computed binary target by treating “directional match” as the positive class (i.e., TP: signs match; FP: signs match when the reference does not; FN: signs do not match when the reference does). In practice, this means micro-averaged recall is always equal to the MDA, since both reduce to the proportion of correctly classified instances; in addition, this reduces each update to one of three labels: Up ($\Delta>\epsilon$), Down ($\Delta<-\epsilon$), or Still ($|\Delta|<\epsilon$); and we evaluate whether the model’s label matches the reference.
@@
To assess whether models produce well-calibrated confidence estimates, we compute the Brier Score (Brier, 1950). This allows us to evaluate whether the model’s confidence aligns with reference before and after new information is introduced. The Brier score between model prediction $p$ and reference forecast $h$ is defined as: $\text{Brier}(p,h)=(p-h)^{2}.$ We report the change in calibration directly as the average difference in squared errors before and after conditioning on new information:
@@
A negative $\Delta\text{Brier}$ indicates improved calibration after observing new information (since lower Brier is better), while a positive value suggests degradation.
@@ 4 EvolveCast Dataset Construction
EvolveCast is constructed from Metaculus (www.metaculus.com; see examples in Appendix C), an online forecasting platform where users submit probability estimates to questions spanning domains such as politics, economics, health, and technology. Metaculus aggregates these into a community prediction that is continuously updated until shortly before resolution. Each question is associated with predefined resolution criteria, and the platform enforces strict guidelines to ensure that forecasts are consistent with them and human-generated and consistent with those criteria. Metaculus also has a strong empirical track record: over the past five years, community predictions on resolved binary questions have shown close alignment to observed outcomes, with roughly half of observed frequencies lying within the 90% credible interval around the ideal calibration line. [^1] https://www.metaculus.com/questions/track-record This calibration performance provides confidence that the platform’s aggregated predictions are reliable.
We apply several filtering steps to ensure data quality. First, we discard questions that Metaculus marks as ambiguous or that are resolved in ways that make forecasting infeasible. For example, when the outcome falls outside the prediction range or when resolution criteria are modified after submission. Second, we only include questions with at least 100 individual forecasts to maintain statistical reliability in the aggregated reference prediction. Finally, we restrict the dataset to binary (Yes/No) questions. The resulting dataset consists of 1,613 question–news pairs with timestamp.
@@ 4.1 Reference Forecasts and Confidence Extraction
In our setting, ground-truth confidence values are not directly observable. We therefore follow prior work that uses human forecast distributions as a proxy for uncertainty (Plank, 2022; Baan et al., 2024; Yuan et al., 2025). The intuition is that aggregated human forecasts, even when forecasters disagree, provide informative signals of uncertainty: disagreement itself reflects the inherent ambiguity and difficulty of the task, similar to how disagreement among annotators is leveraged in natural language tasks with multiple plausible interpretations. This approach has several advantages. First, it captures the fact that some forecasting questions are inherently uncertain, so variance across forecasters represents a genuine property of the task rather than annotation noise. Second, aggregation over many individuals smooths out idiosyncratic biases while preserving collective uncertainty, yielding a stable yet informative signal. Third, these distributions are updated dynamically as new evidence appears, making them particularly well suited for evaluating belief revision.
For Boolean questions, where the community assigns probability $P$ to the correct outcome, we compute a normalized confidence score as $h=\sigma\!\left(\frac{\ln P-\ln 0.5}{\ln 2}\right).$ This transformation ensures that confidence is scaled relative to a chance-level (50%) baseline. While aggregated forecasts are sometimes incorrect, they nonetheless represent the best available reasoning given the information at the time, and thus serve as a practical and informative reference point for evaluating model confidence.
@@ 4.2 News Retrieval and Alignment
To select the most relevant update, we compute semantic similarity between the comment text and each retrieved article using sentence embeddings (Reimers & Gurevych, 2019). Articles are then ranked by similarity score, and by default we retain the top-ranked article, storing both its title and headline as the associated news update. This procedure ensures that each question–forecast pair is aligned with a news item that is both temporally plausible and semantically linked to the corresponding forecaster discussion. Because a single question may be associated with multiple updates over time, the dataset may contain repeated questions paired with different news events. In total, this process yields 1,613 aligned question–news pairs across domains.

### v2

# Do Language Models Update their Forecasts with New Information?

@@ Abstract
Prior work has largely treated forecasting as a static task, failing to consider how forecasts and the confidence in them should evolve as new evidence emerges. To address this gap, we introduce EvolveCast, a framework for evaluating whether large language models revise their forecasts appropriately in response to new information. In particular, EvolveCast assesses whether LLMs update their forecasts when presented with information released after their training cutoff. We use human forecasters as a comparative reference to assess forecast updates and confidence calibration under new information. While LLMs demonstrate some responsiveness to new information, their updates are often inconsistent or overly conservative. We further find that both verbalized and logits-based confidence estimates remain far from the human reference standard. Across settings with a variety of LLMs, models tend to be conservative in updating their forecasts. These findings suggest that current approaches (e.g., RAG-based methods) for “updating” model knowledge are insufficient for probabilistic reasoning; models treat new information as retrieval context rather than evidence that shifts posterior probability. EvolveCast thus underscores the need for more robust mechanisms to incorporate external knowledge into belief dynamics.
Machine Learning, ICML
@@ 1 Introduction
Large language models have achieved strong performance across a wide range of NLP tasks, particularly those involving factual recall and reasoning over known information (Lin et al., 2022; Guo et al., 2023). However, most evaluations remain static and retrospective: they assess what models know about the world up to their training cutoff. In contrast, real-world decision-making often requires reasoning about uncertain future events, where outcomes are not yet known and relevant information evolves over time.
Forecasting plays a central role in domains such as policymaking, science, and technology. It requires anticipating future developments based on incomplete or uncertain evidence, and adjusting those beliefs as new signals arrive. Unlike fact retrieval tasks, forecasting is inherently temporal and dynamic. For instance, given the question, “Will GPT-6 be released in 2026?”, an LLM might answer “Yes” based on existing trends in AI development. However, without a way to incorporate emerging information, such as announcements or delays, such forecasts remain static, and may quickly become outdated or misleading. To illustrate the dynamic nature of forecasting, see the example in Fig 1.
QuestionModelReference Q: Will GPT-6 be released in 2026? Initial forecast (t=0t{=}0)
Prediction:
Yes    Conf: $55\%$
News update: “Sam Altman hints at major 2026 release.” Updated forecast (t=1t{=}1)
Prediction:
Yes
Confidence: $60\%$ (Up $5\%$)
@@
Optional Input(History)Model under-reacts vs. reference (1010 $p$ shortfall)
Figure 1: The EvolveCast framework. Models are evaluated on their ability to update forecasts ($P_{0}\rightarrow P_{t}$) when exposed to new evidence ($x_{t}$). The dashed arrow indicates an optional setting (Sec. 6.4) where the model leverages the historical human reference ($r_{0}$) as context. While the model correctly interprets the news as a positive signal (Direction: Up), it revises its confidence conservatively compared to the human reference, illustrating a magnitude mismatch (Up $5\%$ vs. Up $15\%$).
Prior work has begun to explore the forecasting capabilities of LLMs (Jin et al., 2020; Zou et al., 2022; Yuan et al., 2024a). These efforts typically frame forecasting as a static problem, where the model produces a categorical answer based on its parametric knowledge. Some studies incorporate external information, including textual evidence such as news (Ye et al., 2024; Halawi et al., 2024), but use this information to improve the model’s final prediction, not to evaluate how forecasts change over time. More recently, forecasting has been reframed as a probabilistic task, where models are expected to produce confidence scores in addition to predictions (Karger et al., 2024; Yuan et al., 2025). In this setting, the goal is not just to answer the mentioned question with “Yes”, but to provide a confidence level, e.g., 70%, that reflects the model’s uncertainty. These works though showed that even when models make correct predictions, their confidence is often miscalibrated.
Yet even confidence-aware predictions remain static if models cannot revise their beliefs. In practice, forecasters routinely update their views as new information becomes available. For example, if a forecaster initially assigns a 55% chance to the release of GPT-6 in 2025, but later encounters a credible report stating “Sam Altman has hinted at a major announcement next week,” they might revise their estimate to 70%. A capable forecasting model should exhibit similar adaptive behavior, not only producing reasonable confidence scores, but updating them in response to emerging signals. Evaluating belief dynamics provides insight into how well models reason under uncertainty and adapt over time, which remains a blind spot in most LLMs.
This paper introduces EvolveCast, a framework for evaluating whether language models revise their forecasts and confidence in response to emerging new information or evidence. Crucially, we assess the reasoning process of confidence update rather than the static accuracy of the final outcome. Unlike benchmarks measuring performance against a final resolution e.g., whether an event happened or not, Unlike benchmarks measuring performance against a final resolution (e.g., whether an event eventually occurred), we measure alignment with the aggregate human forecast. We treat this aggregation not as ground truth, but as a robust proxy for rational belief updating under uncertainty.
Recall the GPT-6 example: if a credible rumor causes human forecasters to raise their probability from 55% to only 70%, a well-calibrated model should exhibit a similar, conservative shift. Evaluating against the final outcome (which is binary) fails to capture this nuance; a model that guesses 100% might eventually be “correct” about the result, but it would have been hallucinating certainty given the evidence available at the time. This approach distinguishes rational updating, adjusting credence in proportion to evidence, from clairvoyance, i.e., predicting the final outcome before sufficient evidence exists. Revisiting our example: suppose GPT-6 is eventually released in 2025. A metric based on final resolution would favor a model that blindly guessed 100% probability early on, while penalizing a rational model that maintained 55% confidence based on the limited available news. By evaluating against the human aggregate rather than the future outcome, we avoid such erroneous penalties for justified uncertainty. To this end, we construct a benchmark of scenarios where information becomes available after training cutoffs. We also evaluate a setting where models leverage historical human forecasts as a contextual reference (visualized as the optional input path in Figure 1), mimicking real-world scenarios where forecasters observe market consensus.
@@
Our evaluation focuses on the direction and magnitude of prediction shifts and the calibration of updated confidence scores as exampled in Fig 1. Our findings show that while language models can update their forecasts in response to new information, their confidence updates are conservative and confidence estimates vary by method without a clear advantage, and both fall short of human reference forecasts. These challenges highlight fundamental difficulties in modeling belief dynamics and underscore the need for evaluation beyond static forecasting accuracy. EvolveCast provides such a framework, offering a setting for analyzing how models adapt their predictions as the world evolves.
@@ 2 Related Work
A growing body of research has investigated the forecasting capabilities of language models, though most efforts emphasize event prediction rather than belief dynamics or confidence calibration. OpenForecast (Wang et al., 2025) introduces a large-scale benchmark for open-ended multi-step forecasting, but focuses primarily on model accuracy rather than the calibration of confidence estimates. ForecastBench (Karger et al., 2024) frames forecasting as a probabilistic task with evolving predictions over time, but does not explicitly examine how models adjust their beliefs in light of new information. Time-R1 (Liu et al., 2025) presents a reinforcement learning framework to endow smaller LLMs with temporal understanding and future event generation capabilities, showing strong performance on forecasting tasks beyond training cutoff. While our focus also involves forecasting under temporal uncertainty, unlike previous work we evaluate belief revisions as new signals emerge. Beyond forecasting, several benchmarks target reasoning and plausibility assessment. COPA (Roemmele et al., 2011) and HellaSwag (Zellers et al., 2019) evaluate causal and commonsense reasoning via multiple-choice inference tasks. PRobELM (Yuan et al., 2024b) examines models’ ability to rank outcomes by plausibility, drawing on general world knowledge. While these evaluations probe important reasoning capabilities, they do not test how models reason under uncertainty or update beliefs.
@@ 3.1 Task Definition
We formalize EvolveCast as a dynamic forecasting task, where the goal is to evaluate whether a language model revises its forecasts and confidence in response to new information. Let $q$ denote a forecasting question about a future event (e.g., “Will GPT-6 be released in 2025?”). At time $t$, new information $x_{t}$ (e.g., a news headline or update) becomes available, and we measure how the model’s belief changes when conditioned on it. For binary questions ($\mathcal{Y}=\{\text{Yes},\text{No}\}$), the model’s belief at time $t$ reduces to a scalar confidence $p_{t}=P_{t}(\text{Yes}\mid x_{t})\in[0,1]$, and the confidence update is defined as $\Delta p=p_{t}-p_{0}$.
@@ 3.2 Evaluation Criteria
We evaluate the quality of model confidence updates by comparing the change in model confidence, $\Delta p=p_{t}-p_{0}$, against the shift in a reference forecast, $\Delta h=h_{t}-h_{0}$ (the construction of which is detailed in §4). It is important to note that a key methodological choice here is to compare against $\Delta h$ rather than the final resolution ($y\in\{0,1\}$). The final resolution captures the eventual state of the world at a later time but does not represent the justified belief state at time $t$. For example, if an event is 60% likely given current news, the rational prediction is $0.6$; predicting $1.0$ is an error of overconfidence, even if the event eventually occurs. Comparing confidence updates against the final resolution encourages the model to act as an oracle, penalizing valid epistemic uncertainty. By comparing against the reference $h$, we focus on the reasoning process: How should a rational agent’s uncertainty update given this specific piece of news?
We assess this alignment on three aspects:
@@
We use Mean Directional Accuracy (MDA) to assess whether the model and reference forecasts tend to update their beliefs in the same direction:
@@
for a set of $N$ instances where $\Delta p_{i}$ and $\Delta h_{i}$ are the model and reference forecast deltas for the $i$-th instance. We apply a minimal change threshold $\epsilon$ to avoid counting negligible shifts as meaningful updates. Specifically, we treat both $\Delta p_{i}$ and $\Delta h_{i}$ as zero when $|\Delta|<\epsilon$, and consider the signs to match in such cases. Let $\tilde{\Delta}_{p}$ denote the thresholded version of $\Delta p$, where $\tilde{\Delta}_{p}=0$ if $|\Delta p|<\epsilon$, and similarly for $\Delta h$. Based on this thresholding, we map each instance to one of three directional labels: Up ($\tilde{\Delta}>0$), Down ($\tilde{\Delta}<0$), or Still ($\tilde{\Delta}=0$). We then evaluate the model’s ability to predict these reference labels using standard multi-class Precision, Recall, and F1 scores.
@@
To assess whether models produce well-calibrated confidence estimates, we compute the Brier Score (Brier, 1950). This allows us to evaluate whether the model’s confidence aligns with reference before and after new information is introduced. The Brier score between model prediction $p$ and reference forecast $h$ is defined as: $\text{Brier}(p,h)=(p-h)^{2}.$ The change in calibration directly as the average difference in squared errors before and after conditioning on new information:
@@
A negative $\Delta\text{Brier}$ indicates improved calibration after observing new information (lower Brier is better), while a positive value suggests degradation.
@@ 4 EvolveCast Dataset Construction
EvolveCast is constructed from Metaculus (www.metaculus.com; see examples in Appendix C), an online forecasting platform where users submit probability estimates to questions spanning domains such as politics, economics, health, and technology. Metaculus aggregates these into a community prediction that is continuously updated until shortly before resolution. Each question is associated with predefined resolution criteria, and the platform enforces strict guidelines to ensure that forecasts are consistent with those criteria. Metaculus also has a strong empirical track record: over the past five years, community predictions on resolved binary questions have shown close alignment to observed outcomes, with roughly half of observed frequencies lying within the 90% credible interval around the ideal calibration line. [^1] https://www.metaculus.com/questions/track-record This calibration performance provides confidence that the platform’s aggregated predictions are reliable to be used as references.
We apply several filtering steps to ensure data quality. First, we discard questions that Metaculus marks as ambiguous or that are resolved in ways that make forecasting infeasible. For example, when the outcome falls outside the prediction range or when resolution criteria are modified after submission. Second, we only include questions with at least 100 individual forecasts to maintain statistical reliability in the aggregated prediction. Finally, we restrict the dataset to binary (Yes/No) questions. These filtered questions form the basis of our benchmark, which we subsequently populate with relevant news updates (as detailed in Section 4.2).
@@ 4.1 Reference Forecasts and Confidence Extraction
In our setting, ground-truth confidence values are not directly observable. We therefore follow prior work that uses human forecast distributions as a proxy for uncertainty (Plank, 2022; Baan et al., 2024; Yuan et al., 2025). The intuition is that aggregated human forecasts, even when forecasters disagree, provide informative signals of uncertainty that reflect the latent probability of the event at that specific moment in time, which the final binary outcome cannot capture. Disagreement itself reflects the inherent ambiguity and difficulty of the task, similar to how disagreement among annotators is leveraged in tasks with multiple plausible interpretations. This approach has several advantages. First, it captures the fact that some forecasting questions are inherently uncertain, so variance across forecasters represents a genuine property of the task rather than annotation noise. Second, aggregation over many individuals smooths out idiosyncratic biases while preserving collective uncertainty, yielding a stable yet informative signal. Third, these distributions are updated dynamically as new evidence appears, making them suited for evaluating belief revision.
For Boolean questions, where the community assigns probability $P$ to the correct outcome, we compute a normalized confidence score as $h=\sigma\!\left(\frac{\ln P-\ln 0.5}{\ln 2}\right).$ This ensures that confidence is scaled relative to a chance-level (50%) baseline. While aggregated forecasts are sometimes incorrect, they nonetheless represent the best available reasoning given the information at the time, and thus serve as a practical and informative reference point for evaluating model confidence.
@@ 4.2 News Retrieval and Alignment
To select the most relevant update, we compute semantic similarity between the comment text and each retrieved article using sentence embeddings (Reimers and Gurevych, 2019). Articles are then ranked by similarity score, and by default we retain the top-ranked article, storing both its title and headline as the associated news update. This procedure ensures that each question–forecast pair is aligned with a news item that is both temporally plausible and semantically linked to the corresponding forecaster discussion. In total, this process yields 1,613 aligned question–news pairs derived from 203 unique questions. This structure allows us to evaluate not just isolated predictions, but the consistency of belief updating over time for the same event.
To validate the quality of the aligned news updates, we conducted a manual inspection of a random subset ($N=20$) of question-news pairs. We found that in 90% of cases, the retrieved article was directly relevant to the forecasting topic and provided information post-dating the question start time. This confirms that the downstream performance gaps stem primarily from reasoning failures (i.e., how to weigh the evidence) rather than retrieval noise (i.e., irrelevant context).